# Scale-aware Recognition in Satellite Images under Resource Constraints

### v1

@@ Abstract
Recognition of features in satellite imagery (forests, swimming pools, etc.) depends strongly on the spatial scale of the concept and therefore the resolution of the images. This poses two challenges: Which resolution is best suited for recognizing a given concept, and where and when should the costlier higher-resolution (HR) imagery be acquired? We present a novel scheme to address these challenges by introducing three components: (1) A technique to distill knowledge from models trained on HR imagery to recognition models that operate on imagery of lower resolution (LR), (2) a sampling strategy for HR imagery based on model disagreement, and (3) an LLM-based approach for inferring concept “scale". With these components we present a system to efficiently perform scale-aware recognition in satellite imagery, improving accuracy over single-scale inference while following budget constraints. Our novel approach offers up to a 26.3% improvement over entirely HR baselines, using 76.3 % fewer HR images.
@@ 1 Introduction
The ever-increasing number of earth observation satellites (now more than 1500) offer us a unique opportunity to understand changes at the planetary scale, be it tracking the ecological degradation of forests and coral reefs [47, 16], the destruction of cultural heritage [40], or economic development and well-being [13]. A key to all these downstream applications is the ability to recognize accurately a broad vocabulary of concepts: a major computer vision challenge.
An important aspect of recognition in satellite imagery is the notion of scale: Satellite image resolution is dependent on the particular satellite/sensor and is characterized by the GSD, or ground sampling distance: the distance corresponding to a single pixel. For example, Sentinel-2 images are fairly low resolution (1 pixel = 100 m2[14], while images captured by the NAIP program are fairly high resolution (1 pixel = 1 m2[42]. The image scale matters because many concepts of interest also have a characteristic physical size/scale. An olympic-sized swimming pool, typically about 50 m long, would be barely a speck in Sentinel-2 images. In contrast, a lake may be hundreds of square kilometers, necessitating an exceptionally large image to capture its extent in terms of NAIP imagery, which current models would not be equipped to do.
@@
Accuracy notwithstanding, we must also consider cost and availability. Low-resolution (LR) imagery is free, abundant, and covers the planet densely over space and time [14, 43]. In contrast, high resolution (HR) imagery typically comes from drones or low flying satellites that are commissioned as needed, and therefore come with a high cost and inconsistent temporal and spatial coverage [31, 42].
Scientists in many application domains, be it ecology, archaeology, and urban planning [47, 16, 40, 13], today manually reason about these different aspects, namely, the scale of the concept, the nature of the underlying geography, and the cost of acquired imagery, to achieve the best recognition that they can afford with their budget. We need an approach that can automate this tradeoff and automatically identify when to acquire more expensive high resolution imagery, taking into account the cost, the scale of the concept and the geography. While there is past work on the impact of scale on satellite image recognition, much of it is focused on accuracy and disregards costs [34, 26, 3]. Prior work that has looked at costs focuses on the geography but ignores the scale of the concept itself [44]. A holistic treatment of scale to achieve accurate recognition under a budget is missing.
@@ 2.1 Multi-scale Recognition in Satellite Imagery
Remotely sensed images or satellite images are inherently multi-scale. Depending on the sensor heights and properties, the physical distance between two adjacent pixels in remotely sensed images (known as Ground Sample Distance/GSD) could vary from 0.3m to 1km. These multi-scale images provide complementary information for various applications [47, 16, 40, 13].
Few prior works have explicitly investigated the effect of different scales in visual recognition in satellite imagery. For instance, [34] shows that it is crucial to build scale-specific representations for images with different GSD. However, they do not explicitly address the high acquisition cost of high-resolution satellite imagery. Works on image super-resolution [37, 21, 48] seek to sidestep the problem of acquisition cost by upsampling low-resolution to their high-resolution counterparts. But as shown in [48], these models are prone to hallucination and would produce high-resolution imagery with limited fidelity. In this work, we consider a more realistic setup where we allow a small budget for acquiring high-resolution images during inference. This allows us to tap into the stronger performance of scale-aware models such as [34, 26, 3] while avoiding the hallucination problem in image super-resolution.
@@ 2.2 Resource Constraints in Visual Recognition
Various resource constraints could be considered during the development or deployment of visual recognition models. Active learning approaches [41] consider a budget for acquiring annotations during development and seek to develop the best-performing models by selecting the most informative set of training examples to annotate. This could be done by selecting examples that the model is most uncertain about [20, 36, 50], examples that are most representative of the unlabeled data [2, 35], or a combination of both [49]. Prior works have also considered compute constraints for developing and deploying visual recognition models. These works satisfy the compute constraints either through developing specialized model architectures [38] or compressing/pruning existing models [19, 5]. Our work considers a different type of constraint inherent to remote sensing, i.e., the acquisition cost for high-resolution images. While others have looked into optimizing high-resolution image acquisition [44, 27], we provide a solution to produce the performant visual recognition models under a fixed acquisition budget with a semantic understanding of the concept. To the best of our knowledge, this is an uncharted problem domain.
@@ 2.3 Knowledge Distillation/Learning with Privileged Information
To recognize fine-grained concepts in LR images, we construct a query model by distilling the logits generated by another model that is trained on HR imagery. This approach bears a resemblance to knowledge distillation [19, 6, 10, 29, 15, 7], which seeks to compress a large-scale teacher model by training a compute-efficient student model to mimic the output of the teacher model. Different from conventional knowledge distillation, our query student model and teacher model operate on different inputs. Specifically, our query model operates on LR imagery, whereas our teacher model operates on HR imagery. Our approach can also be viewed as an instance of learning with privileged information [46, 30, 45], in which additional information that is available during test time (e.g., high-resolution satellite imagery) is used for training a model. This paradigm has enabled the development of better-performing models in various applications, such as autonomous driving [8], image super-resolution [23], and so on. In remote sensing, prior works [22, 24] have attempted to use additional sensor data as privileged information for improving land-cover classification. Our work leverages high-resolution satellite image data as privileged information for improving recognition in low-resolution imagery: a different scenario.
@@ 3.1 Problem Setup
To allow for algorithms to train any models or validate hyperparameters, we assume access to a fixed set of locations $\mathcal{I}^{train}$ where we have already acquired (a) images from both resolutions, and (b) annotations for the presence/absence of a set of “seen" concepts $\mathcal{C}^{seen}$. For fairness, these locations are distinct from the locations used for evaluation. Our set of evaluation concepts are also kept distinct (“unseen" classes); however, we also report performance on the seen classes. A full list of the concepts, seen and unseen, is in Appendix 6.1.
@@ 3.2 System Overview
3.3 Inferring concept “scale" using LLMs
@@
We address this knowledge gap with in-context learning [12]. Concretely, we first use the available training locations $\mathcal{I}^{train}$ and the annotations therein to evaluate both the HR scores and the LR scores for each concept in $\mathcal{C}^{seen}$. Thus, for each of these seen concepts, we know which modality yields higher accuracy. We then give these seen concepts, as well as the corresponding modality, as in-context examples to the LLM and ask the LLM to infer the right modality for all other (unseen) concepts. The prompts to the LLM can be found in sec. 6.2 in the appendix.
@@ 3.4 Knowledge Distillation from HR to LR
By minimizing the loss, we essentially distill the HR knowledge [19] in the $M^{HR}$ into a model that consumes LR images $M^{LR}_{KD}$. This formulation also bears a resemblance to learning with privileged information [45]. In this case, when training the model $M^{LR}_{KD}$, we use privileged HR information to guide the $M^{LR}_{KD}$ to detect smaller concepts at LR imagery.
@@ 3.6 Scale-Aware Recognition under Resource Constraints
Given a concept $c$, our system uses an LLM-based approach to predict which modality would yield better accuracy irrespective of cost (i.e, whether HR score, $s_{c}^{HR}$, or LR score, $s_{c}^{LR}$ , is better) (Section 3.3). For a concept best suited to LR imagery, no HR imagery is acquired and inference is carried out using the LR model. For concepts for which HR scores are predicted to be better, our system further decides which locations to acquire HR imagery for. We prompt the user for a budget for HR imagery and calculate our disagreement criterion as described in eq. 5. With this criterion and the budget, we select those locations with the maximum disagreement for HR imagery. We then perform inference with the HR model on the acquired HR images. All other locations are scored using LR imagery and the improved KD model, $M^{LR}_{KD}$. The end result is accurate retrieval where HR imagery is used sparingly, only when it is really needed.
@@ 4.1.1 Pre-trained Models
  • GRAFT is a zero-shot vision-language model that can be used for performing open-vocabulary recognition on satellite imagery [26], similar to CLIP [33]. We use the released model for Sentinel-2 and NAIP and train our own model for the NICFI basemaps. The model architecture is ViT-B-16.
@@ 4.1.2 Benchmark
In this benchmark, we use HR imagery (GSD=1m) captured by the National Agriculture Imagery Program [42]. This program only captures data for the US, so this benchmark is restricted to the US. Here, one Sentinel-2 image corresponds to 100 NAIP images. The cost of acquiring similar resolution imagery ranges from \$1.00–\$6.00 per square km [17].
@@
Our second benchmark leverages the NICFI Satellite Data Program Basemaps for Tropical Forest Monitoring [32] which provides HR imagery (GSD=5m) for regions of Central/South America, Africa, and Asia. Similar imagery has prices ranging from \$0.60–\$2.0 per square km [17]. Here one Sentinel-2 image corresponds to 4 NICFI images.
All our data was downloaded via Google Earth Engine [18]. To handle time discrepancies we retrieved the LR image for a set of HR within the same month that the HR image was collected to avoid major disagreements between the two sets of images. This data will be released publicly along with the download scripts so that others may download data for additional locations. For both benchmarks, we follow [26], and use [28] to obtain ground truth annotations for 40 concepts (listed in Appendix 6.1).
@@ 4.1.3 Metrics
Since the scope of this method is retrieval, we evaluate our results with mAP@k averaged over all classes. We also look at the amount of HR imagery utilized, as well as runtime in seconds to perform inference on the imagery. The runtime is calculated on the images determined by each technique; for LR models our test set has 5,015 images, and for our HR models the test set has 501,500 images. Techniques utilizing HR sampling methods, such as our own or Patchdrop [44] use a combination of these two datasets. All inference is performed using a batch size of 32 on a single Nvidia RTX A6000 GPU. For our experiments we set a budget of 1000 locations to acquire HR imagery, with each location covering roughly 5 sq. km. We report our results in sq. km., so we can compare to other techniques that do not sample by location.
@@ 4.2.1 Overall System Performance
  • Patchdrop: Using Patchdrop [44] to determine HR sample locations.
@@
  • nl. sampling: Using sampling strategies put forth in IS-count [27] (via-night lights) and compare it to our own model disagreement technique.
@@
RemoteCLIP liu2024remoteclip [25] LR 0.461 0.493 0.413 0.440 0 17
CLIP-RSICD arutiunian2021fine [1] LR 0.393 0.441 0.311 0.303 0 17
OpenCLIP cherti2023reproducible [9] LR 0.481 0.505 0.392 0.447 0 16
GRAFT LR + nl. sampling meng2022count [27] LR,HR 0.379 0.354 0.440 0.492 5,954 323
@@
SatMAE cong-22 [11] LR 0.351 0.388 0 67
Cross-Scale MAE tang2023cross [39] LR 0.412 0.422 0 91
Supervised LR + nl. sampling meng2022count [27] LR,HR 0.463 0.480 5,954 319
Supervised + Patchdrop uzkent2020learning [44] LR,HR 0.445 0.470 6,776 360
@@
4.2.2 Detection in Low Resolution
We compare the performance of several models for multi-label classification in Tables 3 and 4 on 40 classes (as listed in the appendix 6.1) using LR imagery alone. We see improvements in performance through our knowledge distillation process for both zero-shot and fully-supervised techniques. For each corresponding training data type and model architecture, we also show the performance of the HR model (in gray).
We also compare our zero-shot KD Model with other zero-shot baselines CLIP-RSICD [1] and RemoteCLIP [25] in Table 3. We compare our fully supervised KD Models with other fully supervised baselines ScaleMAE [34], SatMAE [11], and Cross-Scale MAE [39] as well as a Resnet-50 fine-tuned on imagery of multiple resolutions (Multi-Res) in Table 4.

### v2

@@ Abstract
Recognition of features in satellite imagery (forests, swimming pools, etc.) depends strongly on the spatial scale of the concept and therefore the resolution of the images. This poses two challenges: Which resolution is best suited for recognizing a given concept, and where and when should the costlier higher-resolution (HR) imagery be acquired? We present a novel scheme to address these challenges by introducing three components: (1) A technique to distill knowledge from models trained on HR imagery to recognition models that operate on imagery of lower resolution (LR), (2) a sampling strategy for HR imagery based on model disagreement, and (3) an LLM-based approach for inferring concept “scale”. With these components we present a system to efficiently perform scale-aware recognition in satellite imagery, improving accuracy over single-scale inference while following budget constraints. Our novel approach offers up to a 26.3% improvement over entirely HR baselines, using 76.3 % fewer HR images.
@@ 1 Introduction
The ever-increasing number of earth observation satellites (now more than 1500) offer us a unique opportunity to understand changes at the planetary scale, be it tracking the ecological degradation of forests and coral reefs (Wicaksono et al., 2021; Gao et al., 2020), the destruction of cultural heritage (Tapete et al., 2021), or economic development and well-being (Engstrom et al., 2022). A key to all these downstream applications is the ability to recognize accurately a broad vocabulary of concepts: a major computer vision challenge.
An important aspect of recognition in satellite imagery is the notion of scale: Satellite image resolution is dependent on the particular satellite/sensor and is characterized by the GSD, or ground sampling distance: the distance corresponding to a single pixel. For example, Sentinel-2 images are fairly low resolution (1 pixel = 100 m2(ESA, 2024), while images captured by the NAIP program are fairly high resolution (1 pixel = 1 m2(U.S. Geological Survey, 2024). The image scale matters because many concepts of interest also have a characteristic physical size/scale. An olympic-sized swimming pool, typically about 50 m long, would be barely a speck in Sentinel-2 images. In contrast, a lake may be hundreds of square kilometers, necessitating an exceptionally large image to capture its extent in terms of NAIP imagery, which current models would not be equipped to do.
@@
Accuracy notwithstanding, we must also consider cost and availability. Low-resolution (LR) imagery is free, abundant, and covers the planet densely over space and time (ESA, 2024; USGS Earth Resources Observation and Science, 2024). In contrast, high resolution (HR) imagery typically comes from drones or low flying satellites that are commissioned as needed, and therefore come with a high cost and inconsistent temporal and spatial coverage (Planet Labs PBC, 2018–; U.S. Geological Survey, 2024).
Scientists in many application domains, be it ecology, archaeology, and urban planning (Wicaksono et al., 2021; Gao et al., 2020; Tapete et al., 2021; Engstrom et al., 2022), today manually reason about these different aspects, namely, the scale of the concept, the nature of the underlying geography, and the cost of acquired imagery, to achieve the best recognition that they can afford with their budget. We need an approach that can automate this tradeoff and automatically identify when to acquire more expensive high resolution imagery, taking into account the cost, the scale of the concept and the geography. While there is past work on the impact of scale on satellite image recognition, much of it is focused on accuracy and disregards costs (Reed et al., 2023; Mall et al., 2024; Bastani et al., 2023). Prior work that has looked at costs focuses on the geography but ignores the scale of the concept itself Uzkent & Ermon (2020). A holistic treatment of scale to achieve accurate recognition under a budget is missing.
@@ 2.1 Multi-scale Recognition in Satellite Imagery
Remotely sensed images or satellite images are inherently multi-scale. Depending on the sensor heights and properties, the physical distance between two adjacent pixels in remotely sensed images (known as Ground Sample Distance/GSD) could vary from 0.3m to 1km. These multi-scale images provide complementary information for various applications (Wicaksono et al., 2021; Gao et al., 2020; Tapete et al., 2021; Engstrom et al., 2022).
Few prior works have explicitly investigated the effect of different scales in visual recognition in satellite imagery. For instance, Reed et al. (2023) shows that it is crucial to build scale-specific representations for images with different GSD. However, they do not explicitly address the high acquisition cost of high-resolution satellite imagery. Works on image super-resolution (Shermeyer & Van Etten, 2019; Kowaleczko et al., 2022; Wolters et al., 2023) seek to sidestep the problem of acquisition cost by upsampling low-resolution to their high-resolution counterparts. But as shown in Wolters et al. (2023), these models are prone to hallucination and would produce high-resolution imagery with limited fidelity. In this work, we consider a more realistic setup where we allow a small budget for acquiring high-resolution images during inference. This allows us to tap into the stronger performance of scale-aware models such as Reed et al. (2023); Mall et al. (2024); Bastani et al. (2023) while avoiding the hallucination problem in image super-resolution.
@@ 2.2 Resource Constraints in Visual Recognition
Various resource constraints could be considered during the development or deployment of visual recognition models. Active learning approaches (Tuia et al., 2011) consider a budget for acquiring annotations during development and seek to develop the best-performing models by selecting the most informative set of training examples to annotate. This could be done by selecting examples that the model is most uncertain about (Houlsby et al., 2011; Shannon, 2001; Yoo & Kweon, 2019), examples that are most representative of the unlabeled data (Ash et al., 2019; Sener & Savarese, 2017), or a combination of both (Yin et al., 2017). Prior works have also considered compute constraints for developing and deploying visual recognition models. These works satisfy the compute constraints either through developing specialized model architectures (Tan & Le, 2019) or compressing/pruning existing models (Hinton et al., 2015; Blalock et al., 2020). Our work considers a different type of constraint inherent to remote sensing, i.e., the acquisition cost for high-resolution images. While others have looked into optimizing high-resolution image acquisition (Uzkent & Ermon, 2020; Meng et al., 2022), we provide a solution to produce the performant visual recognition models under a fixed acquisition budget with a semantic understanding of the concept. To the best of our knowledge, this is an uncharted problem domain.
@@ 2.3 Knowledge Distillation/Learning with Privileged Information
To recognize fine-grained concepts in LR images, we construct a query model by distilling the logits generated by another model that is trained on HR imagery. This approach bears a resemblance to knowledge distillation (Hinton et al., 2015; Borup & Andersen, 2021; Cho & Hariharan, 2019; Park et al., 2019; Fukuda et al., 2017; Borup et al., 2023), which seeks to compress a large-scale teacher model by training a compute-efficient student model to mimic the output of the teacher model. Different from conventional knowledge distillation, our query student model and teacher model operate on different inputs. Specifically, our query model operates on LR imagery, whereas our teacher model operates on HR imagery. Our approach can also be viewed as an instance of learning with privileged information (Vapnik & Vashist, 2009; Pechyony & Vapnik, 2010; Vapnik et al., 2015), in which additional information that is available during test time (e.g., high-resolution satellite imagery) is used for training a model. This paradigm has enabled the development of better-performing models in various applications, such as autonomous driving (Chen et al., 2019), image super-resolution (Lee et al., 2020), and so on. In remote sensing, prior works (Kumar et al., 2021; Li et al., 2022) have attempted to use additional sensor data as privileged information for improving land-cover classification. Our work leverages high-resolution satellite image data as privileged information for improving recognition in low-resolution imagery: a different scenario.
@@ 3.1 Problem Setup
To allow for algorithms to train any models or validate hyperparameters, we assume access to a fixed set of locations $\mathcal{I}^{train}$ where we have already acquired (a) images from both resolutions, and (b) annotations for the presence/absence of a set of “seen” concepts $\mathcal{C}^{seen}$. For fairness, these locations are distinct from the locations used for evaluation. Our set of evaluation concepts are also kept distinct (“unseen” classes); however, we also report performance on the seen classes. A full list of the concepts, seen and unseen, is in Appendix A.1.
@@ 3.2 System Overview
3.3 Inferring concept “scale” using LLMs
@@
We address this knowledge gap with in-context learning (Dong et al., 2022). Concretely, we first use the available training locations $\mathcal{I}^{train}$ and the annotations therein to evaluate both the HR scores and the LR scores for each concept in $\mathcal{C}^{seen}$. Thus, for each of these seen concepts, we know which modality yields higher accuracy. We then give these seen concepts, as well as the corresponding modality, as in-context examples to the LLM and ask the LLM to infer the right modality for all other (unseen) concepts. The prompts to the LLM can be found in sec. A.2 in the appendix.
@@ 3.4 Knowledge Distillation from HR to LR
By minimizing the loss, we essentially distill the HR knowledge (Hinton et al., 2015) in the $M^{HR}$ into a model that consumes LR images $M^{LR}_{KD}$. This formulation also bears a resemblance to learning with privileged information (Vapnik et al., 2015). In this case, when training the model $M^{LR}_{KD}$, we use privileged HR information to guide the $M^{LR}_{KD}$ to detect smaller concepts at LR imagery.
@@ 3.6 Scale-Aware Recognition under Resource Constraints
Given a concept our LLM determines which resolution is best suited (Section 3.3). If the best suited resolution is LR, we perform inference using the LR model and all LR imagery only. If the best suited resolution is HR, we ask the user for their HR budget. Then we perform model disagreement using LR imagery alone and the LR and KD models using eq. 5. With these disagreement scores and the user budget, we then sample the top locations, obtain HR images and perform inference using the HR model for these locations. The retrieval results for the remaining area, that it is preferable to use HR for, but cannot be fit in the budget, is scored using the KD model on LR imagery. The end result is accurate retrieval where HR imagery is used sparingly, only when it is really needed.
@@ 4.1.1 Pre-trained Models
  • GRAFT is a zero-shot vision-language model that can be used for performing open-vocabulary recognition on satellite imagery (Mall et al., 2024), similar to CLIP (Radford et al., 2021). We use the released model for Sentinel-2 and NAIP and train our own model for the NICFI basemaps. The model architecture is ViT-B-16.
@@ 4.1.2 Benchmark
In this benchmark, we use HR imagery (GSD=1m) captured by the National Agriculture Imagery Program (U.S. Geological Survey, 2024). This program only captures data for the US, so this benchmark is restricted to the US. Here, one Sentinel-2 image corresponds to 100 NAIP images. The cost of acquiring similar resolution imagery ranges from \$1.00–\$6.00 per square km (Geocento, 2023).
@@
Our second benchmark leverages the NICFI Satellite Data Program Basemaps for Tropical Forest Monitoring (Planet Team, 2024) which provides HR imagery (GSD=5m) for regions of Central/South America, Africa, and Asia. Similar imagery has prices ranging from \$0.60–\$2.0 per square km (Geocento, 2023). Here one Sentinel-2 image corresponds to 4 NICFI images.
All our data was downloaded via Google Earth Engine (Gorelick et al., 2017). To handle time discrepancies we retrieved the LR image for a set of HR within the same month that the HR image was collected to avoid major disagreements between the two sets of images. This data will be released publicly along with the download scripts so that others may download data for additional locations. For both benchmarks, we follow Mall et al. (2024), and use OpenStreetMap contributors (2024) to obtain ground truth annotations for 40 concepts (listed in Appendix A.1).
@@ 4.1.3 Metrics
Since the scope of this method is retrieval, we evaluate our results with mAP@k averaged over all classes. We also look at the amount of HR imagery utilized, as well as runtime in seconds to perform inference on the imagery. The runtime is calculated on the images determined by each technique; for LR models our test set has 5,015 images, and for our HR models the test set has 501,500 images. Techniques utilizing HR sampling methods, such as our own or Patchdrop (Uzkent & Ermon, 2020) use a combination of these two datasets. All inference is performed using a batch size of 32 on a single Nvidia RTX A6000 GPU. For our experiments we set a budget of 1000 locations to acquire HR imagery, with each location covering roughly 5 sq. km. We report our results in sq. km., so we can compare to other techniques that do not sample by location.
@@ 4.2.1 Overall System Performance
  • Patchdrop: Using Patchdrop (Uzkent & Ermon, 2020) to determine HR sample locations.
@@
  • nl. sampling: Using sampling strategies put forth in IS-count (Meng et al., 2022) (via-night lights) and compare it to our own model disagreement technique.
@@
RemoteCLIP Liu et al. (2024) LR 0.461 0.493 0.413 0.440 0 17
CLIP-RSICD Arutiunian et al. (2021) LR 0.393 0.441 0.311 0.303 0 17
OpenCLIP Cherti et al. (2023) LR 0.481 0.505 0.392 0.447 0 16
GRAFT LR + nl. sampling Meng et al. (2022) LR,HR 0.379 0.354 0.440 0.492 5,954 323
@@
SatMAE Cong et al. (2022) LR 0.351 0.388 0 67
Cross-Scale MAE Tang et al. (2023) LR 0.412 0.422 0 91
Supervised LR + nl. sampling Meng et al. (2022) LR,HR 0.463 0.480 5,954 319
Supervised + Patchdrop Uzkent & Ermon (2020) LR,HR 0.445 0.470 6,776 360
@@
4.2.2 Recognition in Low Resolution
We compare the performance of several models for multi-label classification in Tables 3 and 4 on 40 classes (as listed in the appendix A.1) using LR imagery alone. We see improvements in performance through our knowledge distillation process for both zero-shot and fully-supervised techniques. For each corresponding training data type and model architecture, we also show the performance of the HR model (in gray).
We also compare our zero-shot KD Model with other zero-shot baselines CLIP-RSICD (Arutiunian et al., 2021) and RemoteCLIP (Liu et al., 2024) in Table 3. We compare our fully supervised KD Models with other fully supervised baselines ScaleMAE (Reed et al., 2023), SatMAE (Cong et al., 2022), and Cross-Scale MAE (Tang et al., 2023) as well as a Resnet-50 fine-tuned on imagery of multiple resolutions (Multi-Res) in Table 4.